A study on the latest fad diet claimed that the amounts of weight lost by all people on this diet had a mean of 23.7 pounds and a standard deviation of 7.1 pounds.

Step 1 of 2: if a sampling distribution is created using samples of the amounts of weight lost by 60 people on this diet, what would be the mean of the sampling distribution of sample means? Round to two decimal places, if necessary.

Step 2 of 2: If a sampling distribution is created using samples of the amounts of weight lost by 60 people on this diet, what would be the standard deviation of the sampling distribution of sample means? Round to two decimal places, if necessary.
Answer #1

\\ \mu=23.7, \sigma=7.1\\\\ Step-I\\\\ Mean~of~the~sampling~distribution(\mu_{\bar{x}})=\mu=23.7\\\\ Step-II\\\\ Standard~deviation~of~sample~mean(\sigma_{\bar{x}})=\frac{\sigma}{\sqrt{n}}=\frac{7.1}{\sqrt{60}}=0.9166\simeq 0.92